OpenStudy (gabylovesyou):

What is the equation of the line that passes through the points (-1,3) and (1,11)?

OpenStudy (gabylovesyou):

what is the equation

hero (hero):

Insert each point into y = mx + b creating two equations: 11 = 1m + b 3 =-1m + b Now isolate b in each equation: b = 11 - 1m b = 3 + 1m Add both equations together: 2b = 14 Solve for b: b = 14/2 b = 7 Now find m: m = b - 3 m = 7 - 3 m = 4 y = 4x + 7

OpenStudy (gabylovesyou):

is it 4x - y = -7?

OpenStudy (gabylovesyou):

is it?